# Pathology-Aware Adaptive Watermarking for Text-Driven Medical Image Synthesis

### v1

@@ 1 Introduction
Recent advancements in diffusion probabilistic models [11, 21, 23] have enabled the generation of highly realistic synthetic images, benefiting applications such as content creation [22] and data augmentation [28]. However, as generative models become more powerful, concerns over their potential misuse have grown, leading to the introduction of regulatory measures such as the EU AI Act [7], Korean AI Basic Act [20], and Chinese AI governance rules [27]. In response, deep learning-based watermarking techniques [3, 8, 12, 25, 30, 32] have been developed in the general image domain to help distinguish AI-generated content and prevent unethical use.
While similar text-driven image generative models are now being applied in the medical domain [2, 10, 15, 18, 31], their ability to produce medical images that are well-aligned with the input text raises concerns, as misuse could directly impact clinical decision-making. Any manipulation of medical images can compromise patient care, violate confidentiality, and lead to serious real-world consequences [19]. For instance, fabricated medical images could be exploited for insurance fraud by falsely indicating the presence of a disease. Moreover, synthetic medical images could be altered to create misleading medical records, falsely attributing medical conditions to individuals, potentially leading to legal ramifications and a breach of public trust. Hence, there is an urgent need for watermarking methods that can ensure the authenticity and integrity of text-driven synthetic medical images, safeguarding against potential misuse.
@@ 2.1 Pre-training the watermark extractor
We train the watermark extractor using the HiDDeN architecture [32] as illustrated in Fig. 1(a). This architecture consists of: (1) a watermark encoder $\mathcal{E}_{m}$ embedding a $k$-bit message into an image, (2) a transformation layer, and (3) a watermark extractor $\mathcal{D}_{m}$ that retrieves the binary message. The watermark encoder $\mathcal{E}_{m}$ embeds a $k$-bit message $m\in\{0,1\}^{k}$ into an image $\mathcal{I}$ to produce a watermarked image $\tilde{\mathcal{{I}}}$. After applying image transformation $T$ (e.g., cropping, rotation) to $\tilde{\mathcal{{I}}}$, extractor $\mathcal{D}_{m}$ recovers the message, formally represented as $\tilde{m}=\mathcal{D}_{m}(T(\tilde{\mathcal{{I}}}))$. The model is trained by minimizing the binary cross-entropy loss $\mathcal{L}_{msg}=-\sum_{i=1}^{k}\left[m_{i}\log\sigma(\tilde{m}_{i})+(1-m_{i})\log(1-\sigma(\tilde{m}_{i}))\right]$. Similar to [8], image quality is not optimized (since $\mathcal{E}_{m}$ is discarded); instead, PCA whitening is applied after training to remove bias and decorrelate outputs. Feature vectors from $\mathcal{D}_{m}$ are centered using $\mu=\mathbb{E}[\mathcal{D}_{m}]$, and the covariance matrix $\Sigma=U\Lambda U^{T}$ is eigendecomposed. The whitening transform $\Lambda^{-1/2}U^{T}$ is then applied, with the resulting bias $-\Lambda^{-1/2}U^{T}\mu$ and weight $\Lambda^{-1/2}U^{T}$ appended as a linear layer.
@@ 2.2 Fine-tuning the LDM decoder
To improve the clarity of localization boundaries, we apply an Adaptive Attention Scaling (AAS) mechanism that refines the attention distribution. We first normalize the selected attention map $A_{p}^{*}$ using z-score transformation, followed by smooth sigmoid scaling and global min-max normalization. To emphasize salient regions, we define an adaptive threshold $\tau$ as the 0.7 quantile of $A_{p}^{*}$ and define our final pathology localization map as
@@ 3.1 Evaluation Results
Table 1: Quantitative evaluation of watermarking performance on the MIMIC-CXR-JPG dataset [13] and OIA-ODIR dataset [16]. Evaluated attack scenarios include None (no attack), Brt (brightness ×$\times 2$), Crp (50% center crop), JPG (JPEG compression 50%), Rot (25° rotation), and Res (70% scaling). Method with $\dagger$ uses 32-bit message encoding, while others use 48-bit. We highlight the best and second-best results.
@@
Pna. Pmtx. Atel. Ede. Cmgl. Effu. Avg ()($\downarrow$)( ↓ ) AvgAcc ()($\uparrow$)( ↑ ) PSNR ()($\uparrow$)( ↑ )

### v2

@@ 1 Introduction
Recent advancements in diffusion probabilistic models [11, 12, 22, 24] have enabled the generation of highly realistic synthetic images, benefiting applications such as content creation [23] and data augmentation [29]. However, as generative models become more powerful, concerns over their potential misuse have grown, leading to the introduction of regulatory measures such as the EU AI Act [7], Korean AI Basic Act [21], and Chinese AI governance rules [28].
A widely accepted solution to mitigate such risks is watermarking, which embeds imperceptible yet robust signals into generated content. Traditional signal-based methods [1] are fragile to transformations, prompting the rise of deep learning-based approaches [8, 13, 26, 33] that embed in image or feature space. Recent semantic-aware strategies [3, 31] enhance durability by manipulating latent features, though they may affect image structure. As a result, general-purpose deep watermarking methods continue to gain traction for balancing robustness and fidelity.
This need becomes even more critical in the medical domain, where text-driven image generation [2, 10, 16, 19, 32] enables the synthesis of highly realistic images from clinical descriptions. While effective, such capabilities raise concerns, as misuse could directly impact clinical decisions. Any manipulation of medical images can compromise patient care, violate confidentiality, and lead to serious real-world consequences [20]. For instance, fabricated medical images could be exploited for insurance fraud by falsely indicating the presence of a disease. Moreover, synthetic medical images could be altered to create misleading medical records, falsely attributing medical conditions to individuals, potentially leading to legal ramifications and a breach of public trust. Hence, there is an urgent need for watermarking methods that can ensure the authenticity and integrity of text-driven synthetic medical images, safeguarding against potential misuse.
@@ 2.1 Pre-training the watermark extractor
We train the watermark extractor using the HiDDeN architecture [34] as illustrated in Fig. 1(a). This architecture consists of: (1) a watermark encoder $\mathcal{E}_{m}$ embedding a $k$-bit message into an image, (2) a transformation layer, and (3) a watermark extractor $\mathcal{D}_{m}$ that retrieves the binary message. The watermark encoder $\mathcal{E}_{m}$ embeds a $k$-bit message $m\in\{0,1\}^{k}$ into an image $\mathcal{I}$ to produce a watermarked image $\tilde{\mathcal{{I}}}$. After applying image transformation $T$ (e.g., cropping, rotation) to $\tilde{\mathcal{{I}}}$, extractor $\mathcal{D}_{m}$ recovers the message, formally represented as $)\tilde{m}=\mathcal{D}_{m}(T(\tilde{\mathcal{{I}}}$)). The model is trained by minimizing the binary cross-entropy loss $\mathcal{L}_{msg}=-\sum_{i=1}^{k}\left[m_{i}\log\sigma(\tilde{m}_{i})+(1-m_{i})\log(1-\sigma(\tilde{m}_{i}))\right]$. Similar to [8], image quality is not optimized (since $\mathcal{E}_{m}$ is discarded); instead, PCA whitening is applied after training to remove bias and decorrelate outputs. Feature vectors from $\mathcal{D}_{m}$ are centered using $\mu=\mathbb{E}[\mathcal{D}_{m}]$, and the covariance matrix $\Sigma=U\Lambda U^{T}$ is eigendecomposed. The whitening transform $\Lambda^{-1/2}U^{T}$ is then applied, with the resulting bias $\Lambda^{-1/2}U^{T}\mu$ and weight $\Lambda^{-1/2}U^{T}$ appended as a linear layer.
@@ 2.2 Fine-tuning the LDM decoder
To improve the clarity of localization boundaries, we apply an Adaptive Attention Scaling (AAS) mechanism that refines the attention distribution. We first normalize the selected attention map $A_{p}^{*}$ using z-score transformation, followed by smooth sigmoid scaling and global min-max normalization. To emphasize salient regions, we define an adaptive threshold $\tau$ as the 0.7 quantile of $A_{p}^{*}$, a value empirically determined to provide a favorable trade-off between watermark robustness and the preservation of critical diagnostic information. We then define our final pathology localization map as
@@ 3.1 Evaluation Results
Table 1: Quantitative evaluation of watermarking performance on the MIMIC-CXR-JPG dataset [14] and OIA-ODIR dataset [17]. Evaluated attack scenarios include None (no attack), Brt (brightness $\times 2$), Crp (50% center crop), JPG (JPEG compression 50%), Rot (25° rotation), and Res (70% scaling). Method with $\dagger$ uses 32-bit message encoding, while others use 48-bit. We highlight the best and second-best results.
@@
Pna. Pmtx. Atel. Ede. Cmgl. Effu. Avg $(\downarrow)$ AvgAcc $(\uparrow)$ PSNR $(\uparrow)$
@@ 4 Conclusion and Discussion
4.0.1 Acknowledgements
This work was supported in part by the IITP RS-2024-00457882 (AI Research Hub Project), IITP 2020-II201361, NRF RS-2024-00345806, NRF RS-2023-00219019, and NRF RS-2023-002620.